OpenStudy (anonymous):

FOR FLUCK SAKES, HELP ME: square root negative 10 ( 11+ square root negative 5) A. Negative 5 square root end root plus 11 I square root 10 B. 5 square root 2 end root plus 11 I square root 10 XXXXX C. negative 11 square root of 10 end root plus 5 square root 2 XXXXX XXXXX D. 11 square root of 10 end root plus 5 square root of 2 XXXXX

OpenStudy (anonymous):

use latex for flick sakes :P \[\sqrt{-10}(11 + \sqrt{-5}) = \sqrt{10}i(11 + \sqrt{5}i) = 11\sqrt{10}i - \sqrt{50}\]

OpenStudy (anonymous):

\[ = 11\sqrt{-10} - 5\sqrt{2}\]

OpenStudy (anonymous):

So A? :D

OpenStudy (anonymous):

ya if A as sqrt(-10)

OpenStudy (littlebird):

I have no idea how to correctly read those answer choices.